OpenStudy (anonymous):

Prove that for every positive integer n 1^2 -2^2+3^3-...+(-1)^(n-1) n^2 = (-1)^(n-1) (n(n+1)/2)

OpenStudy (anonymous):

by induction:suppose it is true for n then try to prove it true for n+1

OpenStudy (anonymous):

Okay then I get (-1)^n ((n+1)(n+2))/2

OpenStudy (anonymous):

how do you know it's true though?

OpenStudy (anonymous):

try it first for n=1,2,3...

OpenStudy (anonymous):

okay but how do you show it for any n

OpenStudy (anonymous):

S(n)=1^2 -2^2+3^2-...+(-1)^(n-1) n^2 = (-1)^(n-1) (n(n+1)/2) we want to prove: S(n+1)=(-1)^n ((n+1)(n+2))/2 but: S(n+1)=S(n)+(-1)^(n)( n+1)^2 so S(n+1)=(-1)^(n-1) (n(n+1)/2)+(-1)^(n)( n+1)^2 now factor (-1)^(n-1)*(n+1) out: S(n+1)=(-1)^(n-1)*(n+1)[n/2-n-1] =(-1)^(n-1)*(n+1)(-n-2)/2 and since (-1)^(n-1)=(-1)^n*(-1)^(-1)=-(-1)^n then S(n+1)=(-1)^(n)*(n+1)(n+2)/2 which is what we wanted to prove

OpenStudy (anonymous):

sorry but did you get the s(n) in S(n+1)=S(n)+(-1)^(n)( n+1)^2?

OpenStudy (anonymous):

but where did you**

OpenStudy (anonymous):

each sum is the previous one + the last term((-1)^(n-1) n^2) which depends on n..

OpenStudy (anonymous):

in our case i just put for the last term n+1 instead of each n

OpenStudy (anonymous):

Oh I see thank you that make sense

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

sorry I have one more question about this I just don't see how S(n) which is the sum of all the previous term can be equal to just the previous term (-1)^(n-1) (n(n+1)/2)?

OpenStudy (anonymous):

is (-1)^(n-1) (n(n+1)/2) not just giving us 1 term

OpenStudy (anonymous):

one in my sentence meant previous sum...not term

OpenStudy (anonymous):

OH I SEE

OpenStudy (anonymous):

you took the sum which it was equal to before n+1

OpenStudy (anonymous):

ohhh I see okay okay thanks a lot haha sorry to bug you

OpenStudy (anonymous):

nope